OpenStudy (anonymous):

A polygon has an area of 225 square meters. If the area is tripled, how does each side length change?

Directrix (directrix):

The polygon is not given to be regular nor similar to the polygon which results from the tripled sides.

Directrix (directrix):

If they are, then this theorem would apply: If two solids are similar, the square of the scale factor of the two solids is equal to the ratio of any two corresponding area measurements of the solids.

Directrix (directrix):

That would give the square of the ratio of the sides to be 225/(3*225)

OpenStudy (anonymous):

Hey, thanks so much for explaining :)

Directrix (directrix):

(s/s3)^2 = (225/(3*225) (s/s3)^2 = (1/(3) s/s3 = √(1/3) where s is the side of the original polygon and s3 is the side of the polygon with tripled area. s/s3 = √(3) / 3

Directrix (directrix):

Do you have answer options?

OpenStudy (anonymous):

no

Directrix (directrix):

S3 = √(3) * s) Each original side is multiplied by a factor of √3

Directrix (directrix):

 Here's a way to think about it.